OpenStudy (queenof17):

The slope of the line containing the points (-2, 3) and (-3, 1) is -2/5 1/2 2

OpenStudy (aveline):

(y1-y2)/(x1-x2)=m (3-1)/(-2-(-3))=m (2)/(1)=m 2=m

OpenStudy (queenof17):

thank you!!